# Polynomial extensions of Raimi’s theorem

### v1

@@ Abstract
Raimi’s theorem guarantees the existence of a partition of $\mathbb{N}$ into two parts with an unavoidable intersection property: for any finite coloring of $\mathbb{N}$, some color class intersects both parts infinitely many times, after an appropriate shift (translation). We establish a polynomial extension of this result, proving that such intersections persist under polynomial shifts in any dimension. Given non-constant polynomials $P^{(1)},\dots,P^{(f)}\in\mathbb{Z}[x]$ with positive leading coefficients and $P^{(j)}(0)=0$, we construct a partition of $\mathbb{N}^{k}$ such that for any coloring of $\mathbb{N}^{k}$ with finitely many colors, there exist $x_{0}\in\mathbb{N}$ and a single color class that meets all partition pieces after shifts by $x_{0}+P^{(j)}(h)$ in each of the $k$ coordinate directions, for every $j$ and infinitely many values $h\in\mathbb{N}$. Our proof exploits Weyl’s equidistribution theory, Pontryagin duality, and the structure of polynomial relation lattices. We also prove finite analogues in cyclic groups and $SL_{2}(\mathbb{F}_{q})$, establishing quantitative bounds for Raimi-type intersections in finite settings.
@@ 1 Introduction
Let $r,t,f\in\mathbb{N}$. Let $P^{(1)},\dots,P^{(f)}\in{\mathbb{Z}}[x]$ be non-constant polynomials with the properties that $P^{(1)}(0)=\cdots=P^{(f)}(0)=0$ and the leading coefficients are positive.
@@
Beyond the infinite setting, we can ask whether similar phenomena occur in finite algebraic structures. We answer this affirmatively, establishing finite analogues in both (abelian) cyclic groups and the non-abelian group $SL_{2}(\mathbb{F}_{q})$. These results demonstrate that Raimi-type intersection properties extend to finite algebraic structures, with quantitative bounds replacing infinite intersections.
@@
Let $r,t\in\mathbb{N}$. Let $G$ be a finite cyclic group of order $N$, written additively. There are constants $\alpha=\alpha(r,t)>0$ and $N_{0}=N_{0}(r,t)$ such that the following holds for every $N>N_{0}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $G$ such that for each partition $\bigcup_{j=1}^{t}F_{j}$ of $G$, there exist $m\in\{1,\dots,t\}$ and an element $h\in G$ such that
@@
If the finite cyclic group $G$ is written multiplicatively, then the conclusion becomes $|(F_{m}\cdot h)\cap E_{i}|\geq\alpha|G|$. If $G$ is a finite field of order $q$, where $q$ is a prime power, then we obtain the same conclusion by applying Theorem 1.4 to the multiplicative cyclic group $\mathbb{F}_{q}^{*}=\mathbb{F}_{q}\setminus\{0\}$, which is cyclic of order q1q-1.
@@ 2.2 Preliminary lemmas
4 Cyclic groups – Proof of Theorem 1.4
Since every finite cyclic group is isomorphic to $\mathbb{Z}_{N}$ for some $N$, it is sufficient to prove the theorem for $G=\mathbb{Z}_{N}$.
Proof of Theorem 1.4.
@@
Remark 4.1.
While every finite abelian group $G$ can be decomposed as a direct sum $\mathbb{Z}_{N_{1}}\oplus\cdots\oplus\mathbb{Z}_{N_{f}}$, extending Theorem 1.4 to this setting is not straightforward. The proof of Theorem 1.4 relies crucially on the ordering and interval structure of $\mathbb{Z}_{N}$, which does not have a natural analog in product groups. A natural approach would apply the theorem to each cyclic factor separately, creating a product partition
$E^{(1)}_{i_{1}}\times\cdots\times E^{(f)}_{i_{f}}\quad\text{with }i_{1},\ldots,i_{f}\in\{1,\ldots,r\}$
yielding $r^{f}$ parts, and seek a shift $h=(h^{(1)},\ldots,h^{(f)})$ that works in all coordinates. However, while a large color class $F_{m}$ has large projections $\pi_{i}(F_{m})$ onto each factor by averaging, the set $F_{m}$ itself may lack sufficient product structure to guarantee that $(F_{m}+h)$ intersects all product sets substantially. Thus, a completely different proof technique would be required for general finite abelian groups.

### v2

@@ Abstract
Raimi’s theorem guarantees the existence of a partition of $\mathbb{N}$ into two parts with an unavoidable intersection property: for any finite coloring of $\mathbb{N}$, some color class intersects both parts infinitely many times, after an appropriate shift (translation). We establish a polynomial extension of this result, proving that such intersections persist under polynomial shifts in any dimension. Given non-constant polynomials $P^{(1)},\dots,P^{(f)}\in\mathbb{Z}[x]$ with positive leading coefficients and $P^{(j)}(0)=0$, we construct a partition of $\mathbb{N}^{k}$ such that for any coloring of $\mathbb{N}^{k}$ with finitely many colors, there exist $x_{0}\in\mathbb{N}$ and a single color class that meets all partition pieces after shifts by $x_{0}+P^{(j)}(h)$ in each of the $k$ coordinate directions, for every $j$ and infinitely many values $h\in\mathbb{N}$. Our proof exploits Weyl’s equidistribution theory, Pontryagin duality, and the structure of polynomial relation lattices. We also prove some finite analogues of the above results for abelian groups and $SL_{2}(\mathbb{F}_{q})$.
@@ 1 Introduction
Let $r,t,k,f\in\mathbb{N}$. Let $P^{(1)},\dots,P^{(f)}\in{\mathbb{Z}}[x]$ be non-constant polynomials with the properties that $P^{(1)}(0)=\cdots=P^{(f)}(0)=0$ and the leading coefficients are positive.
@@
Beyond the infinite setting, we can ask whether similar phenomena occur in finite algebraic structures. We answer this affirmatively, establishing finite analogues in both abelian groups and the non-abelian group $SL_{2}(\mathbb{F}_{q})$. These results demonstrate that Raimi-type intersection properties extend to finite algebraic structures, with quantitative bounds replacing infinite intersections.
@@
Let $r,t\in\mathbb{N}$. Let $G$ be a finite cyclic group of order $N$, written additively. There are constants $\alpha=\alpha(r,t)>0$ and $N_{0}=N_{0}(r,t)$ such that the following holds for every $N\geq N_{0}$. There exists a partition $\bigcup_{i=1}^{r}E_{i}$ of $G$ such that for each partition $\bigcup_{j=1}^{t}F_{j}$ of $G$, there exist $m\in\{1,\dots,t\}$ and an element $h\in G$ such that
@@
The same conclusion holds for any finite abelian group $G$ whose exponent is at least $N_{0}$, albeit with a weaker constant $\alpha^{\prime}(r,t)\geq\frac{\alpha(r,t)}{t}.$
If the finite abelian group $G$ is written multiplicatively, then, of course, the conclusion should read as $|(F_{m}\cdot h)\cap E_{i}|\geq\alpha|G|$.
@@ 2.2 Preliminary lemmas
4 Finite abelian groups – Proof of Theorem 1.4
Since every finite cyclic group is isomorphic to $\mathbb{Z}_{N}$ for some $N$, we first prove the theorem for $G=\mathbb{Z}_{N}$.
Proof of Theorem 1.4 for cyclic groups.
@@
Proof of Theorem 1.4 for general abelian groups.
Assume without loss of generality that $G=\mathbb{Z}_{N}\times G^{\prime}$, where $G^{\prime}$ is a finite abelian group and $N\geq N_{0}$. We denote the zero element in $G^{\prime}$ by $0_{G^{\prime}}$. Fix $r,t\in\mathbb{N}$. Let $E_{1},\dots,E_{r}\subset\mathbb{Z}_{N}$ be the partition family given by Theorem 1.4. Define $\widetilde{E}_{i}:=E_{i}\times G^{\prime}$ of $G$.
Then, $\bigcup_{i=1}^{r}\widetilde{E}_{i}$ forms a partition of $G$, and for every partition $\bigcup_{m=1}^{t}F_{m}$ of $G$, there exist an index $m^{\ast}\in\{1,\dots,t\}$ and a shift $h\in G$ such that
$\big|(F_{m^{\ast}}+h)\cap\widetilde{E}_{i}\big|\ \geq\ \frac{\alpha(r,t)}{t}\,|\widetilde{E}_{i}|\qquad\text{for all }i=1,\dots,r,$
where $\alpha(r,t)>0$ is the same constant as in the first part of the theorem.
We now prove the last statement.
For each $x\in\mathbb{Z}_{N}$ and each $m\in\{1,\dots,t\}$, define
$A_{m}(x)=\{y\in G^{\prime}:(x,y)\in F_{m}\}.$
Since $\bigcup_{m=1}^{t}F_{m}$ is a partition of $G$, the sets $\{A_{m}(x)\}_{m=1}^{t}$ form a partition of $G^{\prime}$ for each fixed $x$. In particular,
$\sum_{m=1}^{t}|A_{m}(x)|=|G^{\prime}|\quad\text{for all }x\in\mathbb{Z}_{N}.$
By the pigeonhole principle, for each $x$ there exists at least one index $m_{0}\in\{1,\dots,t\}$ such that
$|A_{m_{0}}(x)|\geq\frac{|G^{\prime}|}{t}.$
For each $x\in\mathbb{Z}_{N}$, we choose one such $m_{0}$ and denote it by $m(x)$. Define
$C_{m}=\{x\in\mathbb{Z}_{N}:m(x)=m\}\quad\text{for }m=1,\dots,t.$
Then $\bigcup_{m=1}^{t}C_{m}$ is another partition of $\mathbb{Z}_{N}$.
Applying Theorem 1.4 to $\mathbb{Z}_{N}$, we obtain an index $m^{\ast}\in\{1,\dots,t\}$ and an element $h_{1}\in\mathbb{Z}_{N}$ such that, for all $i=1,\dots,r$,
$|(C_{m^{\ast}}+h_{1})\cap E_{i}|\geq\alpha(r,t)N.$ (4)
Let $h=(h_{1},0_{G^{\prime}})\in G$. For a fixed $i\in\{1,\dots,r\}$, we compute
$\big|(F_{m^{\ast}}+h)\cap\widetilde{E}_{i}\big|$ $=\sum_{x\in E_{i}}\big|\{y\in G^{\prime}:(x-h_{1},y)\in F_{m^{\ast}}\}\big|.$
Restricting the sum to those $x$ for which $x-h_{1}\in C_{m^{\ast}}$, gives
$\big|(F_{m^{\ast}}+h)\cap\widetilde{E}_{i}\big|\geq\sum_{\begin{subarray}{c}x\in E_{i}\\ x-h_{1}\in C_{m^{\ast}}\end{subarray}}\big|\{y\in G^{\prime}:(x-h_{1},y)\in F_{m^{\ast}}\}\big|.$
For $x$ with $x-h_{1}\in C_{m^{\ast}}$, we have, by the definition of $C_{m^{\ast}}$, that
$\big|\{y\in G^{\prime}:(x-h_{1},y)\in F_{m^{\ast}}\}\big|=|A_{m^{\ast}}(x-h_{1})|\geq\frac{|G^{\prime}|}{t}.$
Therefore,
$\big|(F_{m^{\ast}}+h)\cap\widetilde{E}_{i}\big|\geq\frac{|G^{\prime}|}{t}\cdot\big|E_{i}\cap(C_{m^{\ast}}+h_{1})\big|.$
Moreover, we know from (4) that
$\big|E_{i}\cap(C_{m^{\ast}}+h_{1})\big|\geq\alpha(r,t)N,$
which implies
$\big|(F_{m^{\ast}}+h)\cap\widetilde{E}_{i}\big|\geq\frac{|G^{\prime}|}{t}\cdot\alpha(r,t)N=\frac{\alpha(r,t)}{t}|G|.$
This completes the proof. ∎